OpenStudy (anonymous):

A prop jet can travel 750 miles in the same length of time that it takes a private plane to travel 300 miles. The speed of the prop jet is 60 miles per hour more than twice the speed of the private plane. What is the speed of the prop jet?

OpenStudy (anonymous):

X is the speed of the prop plane. Y is the speed of the private plane. Use these equations : 300/Y = 750/X X - 60 = Y And the substitution property.

OpenStudy (anonymous):

so x= y + 60 ? then plug it in

OpenStudy (anonymous):

yes but if you plug in X - 60 = Y in place of y first, you get x

OpenStudy (anonymous):

okay thanks :)